OpenStudy (anonymous):

simplify : x+7/2x+12 + 6/x^2-36

OpenStudy (anonymous):

aaye soumyadeep is typing!

OpenStudy (anonymous):

\[\frac{ x+7 }{ 2x+12 } + \frac{ 6 }{ x^{2}-36 }\] \[\frac{ x+7 }{ 2(x+6) }+\frac{ 6 }{ (x+6)(x-6) }\] \[\frac{ (x+7)(x-6) + 12}{ 2(x+6)(x-6) }\] \[\frac{ (x+6)(x-5) }{ 2(x-6)(x+6) }\] \[\frac{ x-5 }{ 2(x-6) }\]